OpenStudy (anonymous):

Find S10 for -1 + -7 +-13 + -19 +...

OpenStudy (anonymous):

@SolomonZelman last one, I promise!

OpenStudy (solomonzelman):

I don't mind. Trig is my best class. I had a 98 in a semester exam, and 100 on the final :)

OpenStudy (eric_d):

wow!

OpenStudy (anonymous):

way smarter than me, good thing we have openstudy haha thanks

OpenStudy (solomonzelman):

\(\large\color{red}{ \rm d=-6}\) \(\large\color{red}{ \rm a_{10}=-1+(-6)(10-1)=-1+(-54)=-55}\) \(\large\color{red}{ \rm S_{10}=~½(a_{10}+a_1)10=½(-55+-1)10=(-56)5=?}\)

OpenStudy (anonymous):

-55?

OpenStudy (solomonzelman):

Read my reply please, how are you getting-55 ?

OpenStudy (solomonzelman):

-55 is only A10, but we are at S10 already, no ?

OpenStudy (anonymous):

Im sorry I cant solve it :(

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Just read my reply in red. I promise, it is not so hard :)

OpenStudy (anonymous):

I keep getting -55 @SolomonZelman